OpenStudy (thatgirl21):

Please help, unsure of how to set up. Will give medal & fan! Thomas is earning $50,000 a year as an accountant. He automatically receives a 2 & 1/2 % raise every year. How much is his salary in one year? five years? twenty years? fifty years?

OpenStudy (vuriffy):

50,000*.025 = 1250. First year would be 50,000+1,250 = 51,250.

OpenStudy (thatgirl21):

So I just keep multiplying by 0.25? for the second one would i just do 50,000*0.025(5)

OpenStudy (vuriffy):

I forgot the formulas for this, but you can consistently do this. :o

OpenStudy (vuriffy):

No, it does not work like how you said.

OpenStudy (thatgirl21):

well shoot :(

OpenStudy (vuriffy):

The overall number changes every time.

OpenStudy (vuriffy):

51250 * .025 for second year.

OpenStudy (thatgirl21):

so i just keep doing that until i reach 5?

OpenStudy (vuriffy):

y = a(1 + r)x

OpenStudy (vuriffy):

where: y = value of the variable after x periods (future compounded value) a = initial value of the variable r = compound growth rate x = number of periods

OpenStudy (vuriffy):

y = 50000(1 + 0.025)^5

OpenStudy (vuriffy):

y = 56750.41064 for 5 years. The exponent is the number of periods.

OpenStudy (vuriffy):

Now, just do y = 50000(1 + 0.025)^20 and y = 50000(1 + 0.025)^50

OpenStudy (thatgirl21):

put it in my calculator and got the same thing (: okay!

OpenStudy (vuriffy):

Try ^20 and ^50. This accounts for the year periods.

OpenStudy (thatgirl21):

for 20 i got 81,930.82201

OpenStudy (vuriffy):

81930.8220145, so yes correct.

OpenStudy (thatgirl21):

when i tried to do 50 it did something weird

OpenStudy (vuriffy):

Scientific notation?

OpenStudy (vuriffy):

x10?

OpenStudy (thatgirl21):

it says 7.88860905E-81

OpenStudy (vuriffy):

y = 50000(1 + 0.025)^50 Try again. That was not correct.

OpenStudy (thatgirl21):

okay lol not sure what i did wrong. now i got 171,855.436

OpenStudy (vuriffy):

That is correct, good job.

OpenStudy (thatgirl21):

thank you so much!!

OpenStudy (thatgirl21):

do you happen to know what to do for the same situation, except a percentage decrease? would i just make the ^# negative?

OpenStudy (vuriffy):

Hmm.

OpenStudy (vuriffy):

One second.

OpenStudy (thatgirl21):

like the question is lorenzo and ashton both earn $46,000 a year. lorenzo must have done something to make his boss angry bc he will now get a decrease of 1.75% every year from now on. however, ashton must be a perfect employee bc she will get a 1.75% increase every year from now on. and i'm just supposed to write equations to determine each of their salaries after n years.

OpenStudy (vuriffy):

Well, for Ashton, it would be y = 46000(1+.0175)^n

OpenStudy (thatgirl21):

okay! i figured lorenzo's equation would be y=46000(1-.0175)^n but I'm not really sure

OpenStudy (vuriffy):

There's a different equation for decaying exponential.

OpenStudy (vuriffy):

P(t) = P0e-rt

OpenStudy (vuriffy):

where: P(t) = the amount of some quantity at time t P0 = initial amount at time t = 0 r = the decay rate t = time (number of periods)

OpenStudy (vuriffy):

P(n) = 46000e^-0.0175(n)

OpenStudy (thatgirl21):

so -0.0175(n) would all be an exponent?

OpenStudy (vuriffy):

No, sorry, it shouldn't be negative.

OpenStudy (vuriffy):

P(n) = 46000*0.0175(n)

OpenStudy (vuriffy):

Example: P(2) = 46000*0.0175(2) = 1600. 46000-1600 = 44400.

OpenStudy (thatgirl21):

You're awesome! Thank you so much.

OpenStudy (vuriffy):

Of course, anytime.